Man owns a car lot. He buys cars for $4500 and sells them for $10,000. If he loses $3500 each year his cars sit on his car lot, how many years can he keep the cars before he profits nothing.
Answer by jorel1380(2518)

10000-4500=5500
5500/3500=55/35=11/7 years on the lot before the man starts to lose money..